OpenStudy (anonymous):

Express the complex number in trigonometric form. -2i

OpenStudy (helder_edwin):

given z=a+ib u have to compute \[\large \rho=|z|=\sqrt{a^2+b^2} \] and \[\large \varphi=\text{Arg}(z) \] which is the angle between the positive side of the x-axis and the vector z

OpenStudy (helder_edwin):

then \[\large z=\rho(\cos\varphi+i\sin\varphi) \]

OpenStudy (jdoe0001):

\(\bf -2i\implies 0-2i\qquad \textit{trigonometric form}\implies \color{red}{r[cos(\theta)+isin(\theta)]}\\ \quad \\ \begin{array}{ccllll} a&b\\ 0&-2i \end{array}\\ \quad \\ \implies r =\sqrt{a^2+b^2}\\ \quad \\ tan(\theta)=\cfrac{b}{a}\implies tan^{-1}[tan(\theta)]=tan^{-1}\left(\cfrac{b}{a}\right)\implies \theta=tan^{-1}\left(\cfrac{b}{a}\right)\)